A geometrical version of Hardy’s inequality

Maria Hoffmann-Ostenhof University of Vienna

This is joint work with T. Hoffmann-Ostenhof and A. Laptev. A Hardy’s type inequality is proven for domains in ℝ^{n}, where the distance to the boundary and the volume of the domain is involved. In particular a result is obtained which gives a positive answer to a question asked by Brezis and Marcus.